# Variational formulations of transport phenomena on combinatorial meshes

### v7

@@ 4.2 Topological operations
where $c_{\bullet}$ on the left is a basis cochain on $C_{\bullet}\mathcal{S}$, while $c_{\bullet}$ on the right is a basis cochain on $C_{\bullet}\mathcal{M}$.
@@
Cup products exist for simplicial wilson2007cochain , cubical arnold2012discrete , and polygonal ptackova2017discrete meshes. In all cases they are graded commutative and posses the graded Leibniz rule, but are not associative (even in dimension 11!). However, for simplices the cup product converges to the smooth wedge product [wilson2007cochain, , Theorem 5.4] and cochains coming from de Rham maps are almost associative [wilson2007cochain, , Theorem 5.9]. Cup product can also be defined easily for products of meshes with using the tensor product of non-associative graded commutative differential graded algebras, thus allowing to define it over prisms as well. However, we are unaware of a cup product for a generic polytopal mesh. Moreover, we want to use a notion of topological orthogonality which works well for cubical meshes. For this reason we will use a construction of a quasi-cubical mesh $\mathcal{K}$ (i.e., all its cells have face lattices isomorphic to those of cubes) from a generic mesh $\mathcal{M}$ of simple polytopes, called the Forman subdivision, outlined in detail in [berbatov2022diffusion, , Section 2]. Another reason to use the Forman subdivision is to allow for more complex material properties (conductivities), as discussed in [berbatov2022diffusion, , Section 4]. In this article’s numerical examples we will use only constant conductivities but the intention is to use the discrete formulations (Sections 3.1 and 3.2) with different conductivities of cells of different dimensions as dictated by practical applications.
@@
The above proposition was proved for cubes in [arnold2012discrete, , Theorem 3.2.3] using a lengthy argument. The same proof can be used there since both identities have to be checked on single quasi-cubical cells (we first map cells of $\mathcal{K}$ to cubes). However, a straightforward proof can be done using the fact that cubical meshes with cup products can be defined as tensor products of graded commutative differential graded algebras (the 11D cubes (segments) together with their cup products). The base case (a segment) is trivially checked by brute force.
@@ 5 Combinatorial mesh calculus variational formulations
The exterior calculus formulations in Section 3 can be translated almost literally into CMC formulations. However, we note that the discrete Hodge star operators are not invertible. Therefore, in principle, the discrete strong, primal weak and mixed weak formulations (both transient and steady-state) are not equivalent to one another. Their continuous derivations do not translate literally in the discrete setting. Because our numerical tests show good results, we use literal translations in this work.
@@
$u(t_{0})$ $u_{0}$ $\mathsf{Y}$ Initial condition Equation 3.37
@@ 5.1 Primal weak formulation
$F\colon C^{0}\mathcal{K}\to\mathbb{R},\quad$ $F(w):=(w\smile f)[\mathcal{K}]\qquad$ $[\mathsf{X}\mathsf{T}^{-1}].$ (5.1e)
@@
$\forall w[\mathsf{Y}]\in\mathop{\rm Ker}\nolimits\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},0},\quad$ $B(w,\frac{\partial u}{\partial t})+A_{\mathcal{D}}(w,u)-A_{\mathcal{A}}(w,u)$ $=F(w)-G(w)\qquad$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (5.2a)
@@
$q(t,c_{\bullet})=\begin{cases}(-\star_{1}\widetilde{\kappa}\delta_{0}u+\widetilde{\pi}u\smile v)(t,c_{\bullet}),&c\notin(\Gamma_{\mathbf{N}})_{D-1}\\ g_{\mathbf{N}}(t,c_{\bullet}),&c\in(\Gamma_{\mathbf{N}})_{D-1}\end{cases}.$ (5.3)
@@
$\forall w[\mathsf{Y}]\in\mathop{\rm Ker}\nolimits\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},0},\quad$ $A_{\mathcal{D}}(w,u)-A_{\mathcal{A}}(w,u)$ $=F(w)-G(w)\qquad$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (5.4a)
@@
$q(c_{\bullet})=\begin{cases}(-\star_{1}\widetilde{\kappa}\delta_{0}u+\widetilde{\pi}u\smile v)(c_{\bullet}),&c\notin(\Gamma_{\mathbf{N}})_{D-1}\\ g_{\mathbf{N}}(c_{\bullet}),&c\in(\Gamma_{\mathbf{N}})_{D-1}\end{cases}.$ (5.5)
@@ 5.2 Mixed weak formulation
$G\colon C^{D-1}\mathcal{K}\to\mathbb{R},\;$ $G(r):=(\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},D-1}r\smile g_{\mathbf{D}})[\Gamma_{\mathbf{D}}]\;$ $[\mathsf{Y}],$ (5.6e)
$F\colon C^{D}\mathcal{K}\to\mathbb{R},\;$ $F(\widetilde{w}):=\langle f,\widetilde{w}\rangle_{\mathcal{K},D}\;$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{L}^{-D}],$ (5.6f)
$\mathtt{flow\_rate}\colon C^{0}\mathcal{K}\to\mathbb{R},\;$ $\mathtt{flow\_rate}(w):=\kappa\delta^{\star}_{D}w+\pi\star_{0}w\smile v\;$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}^{-1}].$ (5.6g)
@@ 5.3 Comparison with other methods
DEC is usually presented as a discretisation method based on simplicial meshes and their circumcentric dual, used for producing diagonal Hodge star matrices, where Hodge star operators act between primal $p$-cochains and dual (Dp)(D-p)-cochains. The method does not define the full metric calculus (inner product, adjoint coboundary operator) and cannot be applied to our mixed formulations.
@@ 6 Simulation examples
    since Hodge stars $\star_{p}$ and $\star_{D-p}$ do not cancel (up to a sign $(-1)^{p(D-p)}$) in the discrete case, we would have to keep Hodge stars in the bilinear forms. Going through exterior calculus formulations and translating the resulting final forms allows us to motivate not using Hodge star. Hodge stars are problematic since they increase the stencil and may have non-full rank for some mesh topologies.
@@
    In Section 6.2 we will simulate a problem on a disk with a quadratic potential. We will be using regular meshes (concentric circles intersected by equally spaced rays), but will illustrate the natural handling of curved cells. This is a feature that may not be available in domain discretisation methods for solving continuum problems. For instance, in many finite elements all cells must be flat polytopes since the degrees of freedom must reconstruct piecewise polynomial spaces.
@@ 6.2 Quadratic potential on a 2D disk
The boundary of $M$ is the unit circle $\partial M=\{(x,y)\in\mathbb{R}^{2}\mid x^{2}+y^{2}=1\}$ Let the Dirichlet boundary $\Gamma_{\mathbf{D}}$ be the right half-circle, while and the Neumann boundary $\Gamma_{\mathbf{N}}$ be the left half-circle, i.e.,
@@ 6.4 Linear potential on an irregular 2D mesh
In the first three examples we observed small relative errors for both potentials and flow rates. A common feature of those examples was that in the respective embedded meshes the topologically orthogonal sides were also geometrically orthogonal. This is not the case for the irregular mesh in the last example, suggesting that non-orthogonality of the embedded mesh is the reason for the high relative errors. Moreover, our tests on a problem with a parallelogram grid showed similarly high errors. A compelling reason for this is that the inner product on $1$-cells does not produce optimal Hodge star when topologically orthogonal edges are not geometrically orthogonal. This can be mended potentially by using non-diagonal inner products, i.e., by introducing additional measures for pairs of $p$-cells having a common boundary (p1)(p-1)-cell. Such an inner product will lead to non-local Hodge star (see [wilson2007cochain, , Section 8]) and adjoint coboundary operators but will still be computationally efficient. Indeed, the resulting linear systems for both primal and mixed formulations use only inner products, and can be expressed as sparse linear systems (although elimination of variables in the mixed formulation will not be effective). Regarding post-processing in the primal weak formulation, it can be reduced to solving a linear system with a sparse positive definite matrix, the matrix of the inner product.
@@ 7 Conclusions
Computational impact: The diagonal structure of key matrices in the mixed formulation enables solution strategies unavailable to conventional methods. Our examples show that CMC handles curved geometries and irregular meshes naturally - features that challenge traditional finite elements. While the examples presented use manufactured solutions for verification, they demonstrate the method’s readiness for complex microstructural problems where different cell dimensions require distinct material properties.

### v8

@@ 4.2 Topological operations
where $c_{\bullet}$ on the left is a basis chain on $C_{\bullet}\mathcal{S}$, while $c_{\bullet}$ on the right is a basis chain on $C_{\bullet}\mathcal{M}$.
@@
Cup products exist for simplicial wilson2007cochain , cubical arnold2012discrete , and polygonal ptackova2017discrete meshes. In all cases they are graded-commutative and posses the graded Leibniz rule, but are not associative (even in dimension 11!). However, for simplices the cup product converges to the smooth wedge product [wilson2007cochain, , Theorem 5.4] and cochains coming from de Rham maps are almost associative [wilson2007cochain, , Theorem 5.9]. Cup product can also be defined easily for products of meshes using the tensor product of non-associative graded-commutative differential graded algebras, thus allowing to define it over prisms as well. However, we are unaware of a cup product for a generic polytopal mesh. Moreover, we want to use a notion of topological orthogonality which works well for cubical meshes. For this reason we will use a construction of a quasi-cubical mesh $\mathcal{K}$ (i.e., all its cells have face lattices isomorphic to those of cubes) from a generic mesh $\mathcal{M}$ of simple polytopes, called the Forman subdivision, outlined in detail in [berbatov2022diffusion, , Section 2]. Another reason to use the Forman subdivision is to allow for more complex material properties (conductivities), as discussed in [berbatov2022diffusion, , Section 4]. In this article’s numerical examples we will use only constant conductivities but the intention is to use the discrete formulations (Sections 3.1 and 3.2) with different conductivities of cells of different dimensions as dictated by practical applications.
@@
The above proposition was proved for cubes in [arnold2012discrete, , Theorem 3.2.3] using a lengthy argument. The same proof can be used there since both identities have to be checked on single quasi-cubical cells (we first map cells of $\mathcal{K}$ to cubes). However, a straightforward proof can be done using the fact that cubical meshes with cup products can be defined as tensor products of graded-commutative differential graded algebras (the 11D cubes (segments) together with their cup products). The base case (a segment) is trivially checked by brute force.
@@ 5 Combinatorial mesh calculus variational formulations
The exterior calculus formulations in Section 3 can be translated almost literally into CMC formulations. Because discrete Hodge star operators are generally not invertible, primal and mixed formulations are not equivalent at the discrete level. In the present work, each formulation is therefore treated as a distinct discrete model rather than as an algebraically equivalent representation of the same problem.
@@
Input parameters, initial and boundary conditions are given in Table 6. The unknowns are given in Table 7. $S_{0}$ and $\widetilde{S_{0}}$ are defined as in Equation 3.25 and Equation 3.26 respectively.
@@
$u(t_{0})$ $u_{0}$ $\mathsf{Y}$ Initial condition for potential Equation 3.40
$Q(t_{0})$ $Q_{0}$ $\mathsf{X}$ Initial condition for amount Equation 3.42
@@ 5.1 Primal weak formulation
$F_{1}\colon C^{0}\mathcal{K}\to\mathbb{R},\quad$ $F_{1}(w):=(w\smile f)[\mathcal{K}]\qquad$ $[\mathsf{X}\mathsf{T}^{-1}],$ (5.1e)
$F_{2}\colon C^{0}\mathcal{K}\to\mathbb{R},\quad$ $F_{2}(w):=(\delta_{0}w\smile(\widetilde{S_{0}}\smile v))[\mathcal{K}]\qquad$ $[\mathsf{X}\mathsf{T}^{-1}].$ (5.1f)
@@
$\forall w[\mathsf{Y}]\in\mathop{\rm Ker}\nolimits\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},0},\quad$ $B(w,\frac{\partial u}{\partial t})+A_{\mathcal{D}}(w,u)-A_{\mathcal{A}}(w,u)$ $=F_{1}(w)+F_{2}(w)-G(w)\qquad$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (5.2a)
@@
$q(t,c_{\bullet})=\begin{cases}(-\star_{1}\widetilde{\kappa}\delta_{0}u+(\widetilde{S_{0}}+\widetilde{\pi}u)\smile v)(t,c_{\bullet}),&c\notin(\Gamma_{\mathbf{N}})_{D-1}\\ g_{\mathbf{N}}(t,c_{\bullet}),&c\in(\Gamma_{\mathbf{N}})_{D-1}\end{cases}.$ (5.3)
@@
$\forall w[\mathsf{Y}]\in\mathop{\rm Ker}\nolimits\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},0},\quad$ $A_{\mathcal{D}}(w,u)-A_{\mathcal{A}}(w,u)$ $=F_{1}(w)+F_{2}(w)-G(w)\qquad$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (5.4a)
@@
$q(c_{\bullet})=\begin{cases}(-\star_{1}\widetilde{\kappa}\delta_{0}u+\widetilde{Q_{0}}\smile v)(c_{\bullet}),&c\notin(\Gamma_{\mathbf{N}})_{D-1}\\ g_{\mathbf{N}}(c_{\bullet}),&c\in(\Gamma_{\mathbf{N}})_{D-1}\end{cases}.$ (5.5)
@@ 5.2 Mixed weak formulation
$G_{1}\colon C^{D-1}\mathcal{K}\to\mathbb{R},\;$ $G_{2}(r):=(\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},D-1}r\smile g_{\mathbf{D}})[\Gamma_{\mathbf{D}}]\;$ $[\mathsf{Y}],$ (5.6e)
$G_{2}\colon C^{D-1}\mathcal{K}\to\mathbb{R},\$$G_{2}(r):=\langle r,\kappa^{-1}(\widetilde{S_{0}}\smile v)\rangle_{\mathcal{K},D-1}\$$[\mathsf{Y}],$ (5.6f)
$F\colon C^{D}\mathcal{K}\to\mathbb{R},\;$ $F(\widetilde{w}):=\langle f,\widetilde{w}\rangle_{\mathcal{K},D}\;$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{L}^{-D}],$ (5.6g)
$\mathtt{flow\_rate}\colon C^{0}\mathcal{K}\to C^{D-2}\mathcal{K},\;$ $\mathtt{flow\_rate}(w):=\kappa\delta^{\star}_{D}w+(\widetilde{S_{0}}+\tilde{\pi}w)\smile v\;$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}^{-1}].$ (5.6h)
@@ 5.3 Comparison with other methods
Recent work has clarified the close relationship between discrete exterior calculus and finite element exterior calculus through generalised Whitney forms guzman2025framework . The present framework aligns with this perspective while emphasising a combinatorial organisation of variables and operators, together with a physically grounded interpretation and workflow oriented toward multi-physics coupling on cell complexes.
@@ 6 Simulation examples
    since Hodge stars $\star_{p}$ and $\star_{D-p}$ do not cancel (up to a sign $(-1)^{p(D-p)}$) in the discrete case, we would have to keep Hodge stars in the bilinear forms. Going through exterior calculus formulations and translating the resulting final forms allows us to motivate not using Hodge star. Non-invertibility and rank deficiency of discrete Hodge operators are common features of structure-preserving discretisations hiptmair2001discrete , including FEEC, and should be regarded as intrinsic rather than pathological.
@@
While similar block-diagonal structures arise in hybridised mixed finite element methods ([boffi2013mixed, , Section 7.2], [arnold1985mixed, , Introduction]), in the present framework this structure follows directly from combinatorial inner products, without the introduction of additional hybridisation variables.
@@
    In Section 6.2 we will simulate a problem on a disk with a quadratic potential. We will be using regular meshes (concentric circles intersected by equally spaced rays), but will illustrate the natural handling of curved cells. While classical finite element and isogeometric methods can represent curved geometries through isoparametric or spline-based constructions, the present approach treats curved cells directly at the level of the cell complex, without reference to polynomial reconstruction spaces.
@@ 6.2 Quadratic potential on a 2D disk
The boundary of $M$ is the unit circle $\partial M=\{(x,y)\in\mathbb{R}^{2}\mid x^{2}+y^{2}=1\}$. Let the Dirichlet boundary $\Gamma_{\mathbf{D}}$ be the right half-circle, while and the Neumann boundary $\Gamma_{\mathbf{N}}$ be the left half-circle, i.e.,
@@ 6.4 Linear potential on an irregular 2D mesh
In the first three examples we observed small relative errors for both potentials and flow rates. A common feature of those examples was that in the respective embedded meshes the topologically orthogonal sides were also geometrically orthogonal. This is not the case for the irregular mesh in the last example, suggesting that non-orthogonality of the embedded mesh is the reason for the high relative errors. Moreover, our tests on a problem with a parallelogram grid showed similarly high errors. As with other structure-preserving discretisations, the numerical accuracy of the proposed formulation depends on mesh geometry, with skewed or non-orthogonal meshes leading to reduced performance, as observed in this example. A compelling reason for this is that the inner product on $1$-cells does not produce optimal Hodge star when topologically orthogonal edges are not geometrically orthogonal. This can be mended potentially by using non-diagonal inner products, i.e., by introducing additional measures for pairs of $p$-cells having a common boundary (p1)(p-1)-cell. Such an inner product will lead to non-local Hodge star (see [wilson2007cochain, , Section 8]) and adjoint coboundary operators but will still be computationally efficient. Indeed, the resulting linear systems for both primal and mixed formulations use only inner products, and can be expressed as sparse linear systems (although elimination of variables in the mixed formulation will not be effective). Regarding post-processing in the primal weak formulation, it can be reduced to solving a linear system with a sparse positive definite matrix, the matrix of the inner product.
@@ 7 Conclusions
Computational impact: The diagonal structure of key matrices in the mixed formulation enables efficient solution strategies. Our examples show that CMC handles curved geometries and irregular meshes naturally. While the examples presented use manufactured solutions for verification, they demonstrate the method’s readiness for complex microstructural problems where different cell dimensions require distinct material properties.